OpenStudy (anonymous):

The volume of a cube is 512 cm3. What would change if the volume of the cube were measured in cubic meters

OpenStudy (anonymous):

The volume would change?

OpenStudy (anonymous):

i am really bad at math

OpenStudy (ybarrap):

So this is the inverse of what we had before. Same volume, because changing units does not change that. But we are going from small cm to larger units, meters. So the number would get smaller by .01^3.

OpenStudy (anonymous):

i didnt really know that much either.

OpenStudy (anonymous):

@lilyryan heres a free medal as a gift :)

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} 1m&100cm\\ 1m^2&100^2cm^2\\ 1m^3&100^3cm^3 \end{array}\implies \begin{array}{llll} cm^3&m^3\\ \hline\\ 100^3& 1\\ 512& x \end{array}\implies \cfrac{100^3}{512}=\cfrac{1}{x}\)